Given f(x)=-x-4, find f(a-3)
Answer by tutorcecilia(2152)

f(x)=-x-4 [substitute (a-3) for the x-term and solve as much as possible
f(a-3)=-x-4
f(a-3)=-(a-3)-4
f(a-3)=-a+3-4[multiply through by (-1)
f(a-3)=-a-1
f(a-3)=a+1 [multiply through by (-1)